# noitemsep,topsep=0pt,parsep=0pt,partopsep=0pt

### v1

@@ Abstract
Although data-driven methods have achieved success in 3D human pose estimation, they often suffer from domain gaps and exhibit limited generalization. In contrast, optimization-based methods excel in fine-tuning for specific cases but are generally inferior to data-driven methods in overall performance. We observe that previous optimization-based methods commonly rely on projection constraint, which only ensures alignment in 2D space, potentially leading to the overfitting problem. To address this, we propose an Uncertainty-Aware testing-time Optimization (UAO) framework, which keeps the prior information of pre-trained model and alleviates the overfitting problem using the uncertainty of joints. Specifically, during the training phase, we design an effective 2D-to-3D network for estimating the corresponding 3D pose while quantifying the uncertainty of each 3D joint. For optimization during testing, the proposed optimization framework freezes the pre-trained model and optimizes only a latent state. Projection loss is then employed to ensure the generated poses are well aligned in 2D space for high-quality optimization. Furthermore, we utilize the uncertainty of each joint to determine how much each joint is allowed for optimization. The effectiveness and superiority of the proposed framework are validated through extensive experiments on two challenging datasets: Human3.6M and MPI-INF-3DHP. Notably, our approach outperforms the previous best result by a large margin of 4.5% on Human3.6M. Our source code will be open-sourced.
1 Introduction
Monocular 3D human pose estimation aims to estimate 3D body joints from a single image. This task plays an important role in various applications, such as computer animation [36], human-computer interaction [29], and action recognition [19], but, due to the depth ambiguity, it is still a challenging task that has drawn extensive attention in recent years.
Figure 1: Schematic diagram of the iterative optimization process for human pose estimation. The red pose represents the ground truth, and the blue pose represents our predicted result. The intensity of color corresponds to the number of iterations, with darker colors indicating more iterations. In (a), poses generated by projection constraints are presented. We observe that the projection constraint alone can sometimes mislead the optimization direction, resulting in the optimized 3D pose far from the ground truth (the red pose). In (b), the introduce of uncertainty constraint effectively alleviates this problem.
@@ 1 Introduction
With the development of deep learning, data-driven methods [21, 42, 18, 47, 2] outperform traditional methods and have dominated the field of 3D human pose estimation. During the training process, these methods implicitly acquire knowledge of camera intrinsic parameters and 3D human pose distributions within specific domains. However, these data-driven methods struggle with cross-domain generalization and in-the-wild scenarios [7, 5].
@@
Historically, traditional solutions usually formulate the 3D human pose estimation task as a pure optimization-based problem [24, 26, 33]. Recently, some human mesh recovery methods [14, 8, 15, 12] also adopted optimization strategies to optimize network parameters based on certain constraints. Optimization-based methods offer two significant benefits. First, they prevent the model from generating unrealistic or out-of-distribution poses, which ultimately boosts the model’s robustness. Second, optimization-based approaches offer the flexibility of adaptive adjustments tailored for specific samples, thereby enhancing the model’s generalization capabilities. While optimization-based methods can alleviate domain gaps by estimating 3D poses on a case-by-case basis, their current performance is inferior to data-driven methods. We observe that the previous optimization-based methods [15, 41, 12] typically rely only on the projection constraint. However, due to the inherent depth ambiguity [18], several potential 3D poses could be projected to the same 2D pose. While alignment in 2D space is assured, this imperfect constraint may lead to overfitting during optimization. Additionally, previous optimization-based methods typically employ pre-trained networks to obtain a good initialization, subsequently optimizing the network parameters. Nevertheless, this approach involves a vast number of parameters to be optimized, which may disrupt the prior information embedded in the pre-trained network.
@@
To solve the above problems, we propose an Uncertainty-Aware testing-time Optimization (UAO) framework for 3D human pose estimation. This optimization framework retains the prior knowledge embedded in the pre-trained model and alleviates the overfitting problem using uncertainty of joints. Specifically, we design an effective network for 2D-to-3D pose lifting, and employ a separate decoder to obtain the uncertainty associated with each joint. In this way, our network can simultaneously generates the 3D results and the uncertainty of each joint based on input 2D pose during the training phase. For the testing-time optimization, our target is to make the optimized poses more reliable. Different from previous optimization-based methods that optimize network parameters, the designed optimization strategy freezes the network parameters and optimizes a latent state. This design has two benefits: a) fixing the network parameters can keep the learned pose prior during the optimization process; b) setting the optimization target as a hidden state instead of the network parameter only requires the storage of a small number of parameters, making it more suitable for online inference scenarios. Furthermore, our optimization process incorporates a set of well-designed constraints to guide the optimization direction. For 2D pose alignment, we adopt the camera intrinsic parameters to project the generated 3D pose to 2D space and then minimize the projection loss. Since we do not know the depth information or ground truth, it is hard to get the 3D alignment. We discover that projection constraints may result in overfitting problems, causing the optimized 3D pose to deviate more from the ground truth. To alleviate this, we propose uncertainty constraint that allows 3D joint with higher uncertainty to have a larger degree of optimization, while keeping joint with smaller uncertainty in its original position. As shown in Figure 1, relying solely on projection constraint may inadvertently lead to an incorrect optimization direction, producing worse results. This issue is effectively mitigated when we further introduce uncertainty constraints. With such a unique design, the optimization process is prevented from generating physically implausible poses, leading to more realistic and accurate results.
@@
  • We propose a novel Uncertainty-Aware testing-time Optimization (UAO) framework for 3D human pose estimation, which freezes the pre-trained model to maintain the pose prior and utilize the uncertainty of joints to alleviate the overfitting problem.
  • We design an effective 2D-to-3D lifting network, which can both 3D output while estimating the uncertainty of each joint.
  • For test time optimization, projection constraint is used to ensure that the generated results are aligned in the 2D pose. Furthermore, we design the uncertainty constraint to determine the degree to which each joint can be optimized, which alleviates the overfitting problem.
  • Extensive experiments demonstrate the effectiveness of the proposed UAO framework, and the potential applicability for real-world scenarios.
2 Related Work
@@
Data-driven methods. Data-driven approaches involve training a deep neural network to directly regress 3D human poses from provided 2D poses. Early attempts simply use fully-connected networks (FCNs) to lift the 2D keypoint to 3D space [21]. Since the human skeleton can be naturally represented as a graph, various graph convolutional networks (GCNs) based methods [44, 2, 35, 47, 9] have been proposed. Recently, transformer-based methods [46, 17, 18, 42] have been widely applied to 3D human pose estimation, achieving new breakthroughs. Although these methods can effectively learn pose patterns from datasets, they often exhibit poor generalization ability when faced with unseen data.
@@
Optimization-based methods. During the early stages, researchers employed pure optimization-based methods [24, 26, 33] for 3D human pose estimation. In the era of deep learning, there are also some methods that integrate optimization strategies. We categorize the existing optimization-based methods into two classes. 1) The first class incorporates a regression-based network with optimization strategies, training the entire structure end-to-end, such as SPIN [15] and PyMAF [40]. 2) The second class follows a post-optimization manner, which is more similar to us, using a pre-trained network to obtain a good initialization, including ISO [41], BOA [8], and EFT [12]. However, all these previous post-optimization methods typically optimize the network parameters based on specific constraints, which may disrupt the deep prior knowledge obtained from datasets. As a comparison, we set the optimization variable as the latent state, and keep the model parameters frozen during the iterative optimization process. Besides, pose-related constraints are imperative to ensure the appropriate optimization direction. Previous optimization-based methods simply rely on the projection constraint [15, 41, 12]. But due to the depth ambiguity, this imperfect constraint may lead to the wrong optimization direction when over-optimized. To alleviate this problem, we further utilize the uncertainty constraint to ensure that the generated pose lies within the manifold of 3D poses.
3 Uncertainty Estimation
@@
For 2D-to-3D lifting, data-driven methods regress the 3D human pose $\mathbf{J}^{3D}\in\mathbb{R}^{K\times 3}$ from given 2D pose $\mathbf{J}^{2D}\in\mathbb{R}^{K\times 2}$, where $K$ is the number of skeleton joints. They typically train the model to learn the transformational relationship between 2D pose $\mathbf{J}^{2D}$ and ground truth 3D pose $\mathbf{J}^{3D}$ by minimizing:
$L_{2D}=||\hat{\mathbf{J}}^{3D}-f(\mathbf{J}^{2D})||_{2},$ (1)
@@
To achieve this, we propose GUMLP, primarily consisting of two parts: a GCN layer and a parallel U-shaped multi-layer perception (UMLP) module, as shown in Figure 2. In contrast to the conventional data-driven model [21, 47], which typically employs only one decoder to make direct predictions, we utilize two distinct decoders to predict $\boldsymbol{\mu}$ and $\boldsymbol{s}$, respectively. The GCN layer focuses on local joint relations and introduces the topological priors of the human skeleton. For the second component, the UMLP module has a bottleneck structure [25, 2] instead of the original MLP in the transformer [32]. The input of the UMLP module $\mathbf{X}$ is first processed by a down-projection layer, followed by a middle layer that maintains the same dimension, and ultimately, is fed into an up-projection layer. The simplified formulation can be defined as:
@@
where $\text{MLP}(\cdot)$ consists of a linear layer and a GELU activation [10], LN denotes the Layer Normalization [1]. Here we adopt the parallel UMLP modules along both channel and spatial dimensions to capture multi-scale and multi-level features.
@@
4 Testing-Time Optimization
4.1 Optimization Strategy

### v2

# Uncertainty-Aware Testing-Time Optimization for 3D Human Pose Estimation

@@ Abstract
Although data-driven methods have achieved success in 3D human pose estimation, they often suffer from domain gaps and exhibit limited generalization. In contrast, optimization-based methods excel in fine-tuning for specific cases but are generally inferior to data-driven methods in overall performance. We observe that previous optimization-based methods commonly rely on projection constraint, which only ensures alignment in 2D space, potentially leading to the overfitting problem. To address this, we propose an Uncertainty-Aware testing-time Optimization (UAO) framework, which keeps the prior information of pre-trained model and alleviates the overfitting problem using the uncertainty of joints. Specifically, during the training phase, we design an effective 2D-to-3D network for estimating the corresponding 3D pose while quantifying the uncertainty of each 3D joint. For optimization during testing, the proposed optimization framework freezes the pre-trained model and optimizes only a latent state. Projection loss is then employed to ensure the generated poses are well aligned in 2D space for high-quality optimization. Furthermore, we utilize the uncertainty of each joint to determine how much each joint is allowed for optimization. The effectiveness and superiority of the proposed framework are validated through extensive experiments on challenging datasets: Human3.6M, MPI-INF-3DHP, and 3DPW. Notably, our approach outperforms the previous best result by a large margin of 5.5% on Human3.6M. Code is available at https://github.com/xiu-cs/UAO-Pose3D.
Index Terms:
3D Human Pose Estimation, Testing-time Optimization, Uncertainty Estimation.
I Introduction
@@
Monocular 3D human pose estimation aims to estimate 3D body joints from a single image. This task plays an important role in various applications, such as computer animation [1], human-computer interaction [2], and action recognition [3]. However, due to the depth ambiguity, it remains a challenging task that has drawn extensive attention in recent years. With the development of deep learning, data-driven methods [4, 5, 6, 7, 8] have dominated the field of 3D human pose estimation. During the training process, these methods implicitly acquire knowledge of camera intrinsic parameters and 3D human pose distributions within specific domains. However, these data-driven methods struggle with cross-domain generalization and in-the-wild scenarios [9, 10].
@@
Historically, traditional solutions usually formulate the 3D human pose estimation task as a pure optimization-based problem [11, 12, 13]. These optimization-based methods offer two significant benefits. First, they prevent the model from generating unrealistic or out-of-distribution poses, which ultimately boosts the model’s robustness. Second, they provide the flexibility of adaptive adjustments tailored for specific samples, thereby enhancing the model’s generalization capabilities. Although traditional methods can alleviate domain gaps by estimating 3D poses on a case-by-case basis, their performance is often compromised due to the limitations of manually crafted feature extractors [14]. In contrast, data-driven methods leverage neural networks to learn feature representations in an automated manner, enabling them to effectively capture the intricacy and variability inherent in the data. Recently, some human mesh recovery methods [15, 16, 17, 18, 19] have also adopted optimization strategies to refine network parameters based on certain constraints. We observe that the previous optimization-based methods [18, 20, 19] typically rely only on the projection constraint. However, due to the inherent depth ambiguity [6], several potential 3D poses could be projected to the same 2D pose. While alignment in 2D space is assured, this imperfect constraint may lead to overfitting during optimization. Additionally, previous optimization-based methods typically employ pre-trained networks to obtain a good initialization, subsequently optimizing the network parameters. Nevertheless, this approach involves a vast number of parameters to be optimized, which may disrupt the prior information embedded in the pre-trained network.
@@
Based on the above observations, we propose an Uncertainty-Aware testing-time Optimization (UAO) framework for 3D human pose estimation. This optimization framework retains the prior knowledge embedded in the pre-trained model and alleviates the overfitting problem using the uncertainty of joints. Specifically, we design an effective network for 2D-to-3D pose lifting, and employ a separate decoder to obtain the uncertainty associated with each joint. In this way, our network can simultaneously generate the 3D results and the uncertainty of each joint based on input 2D pose during the training phase. For the testing-time optimization, our target is to make the optimized poses more reliable. Different from previous optimization-based methods [20, 17, 16, 19, 21, 22] that update the network parameters based on specific constraints during test time, our designed optimization strategy freezes the network parameters and optimizes a latent state. This design has two benefits: a) fixing the network parameters can keep the learned pose prior during the optimization process; b) setting the optimization target as a hidden state instead of the network parameter only requires the storage of a small number of parameters, making it more suitable for online inference scenarios. Furthermore, our optimization process incorporates a set of well-designed constraints to guide the optimization direction. For 2D pose alignment, we adopt the camera intrinsic parameters to project the generated 3D pose onto the 2D space and minimize the projection loss. Since we do not know the depth information, achieving precise 3D alignment remains challenging. We observe that relying solely on projection constraint can lead to over-optimization problems, causing the optimized 3D pose to deviate further from the ground truth. To alleviate this, we introduce an uncertainty constraint that allows 3D joint with higher uncertainty to have a larger divergence radius relative to its initial 3D position, while keeping joint with smaller uncertainty closer to its initial 3D position. As shown in Figure 1, relying solely on projection constraint may inadvertently steer optimization in the wrong direction, leading to poorer results. This issue is effectively mitigated by further incorporating uncertainty constraints. With such a unique design, the optimization process is prevented from generating physically implausible poses, leading to more realistic and accurate results.
Figure 1: Schematic diagram of the iterative optimization process. The red pose represents the ground truth, and the blue pose represents the prediction; darker colors indicate later iterations. ‘GT’ refers to the ground truth 3D pose distribution, ‘Init’ denotes the initial result, ‘P’ represents the projection constraint, and ‘P+U’ combines projection and uncertainty constraints. (a) Using only the projection constraint during test-time optimization can lead to significant deviations from the ground truth. (b) Incorporating the uncertainty constraint effectively mitigates this issue, yielding final poses that more closely match the ground truth.
@@
  • We propose a novel Uncertainty-Aware testing-time Optimization (UAO) framework for 3D human pose estimation. First, an effective 2D-to-3D lifting network is designed to generate 3D outputs while estimating the joint uncertainties. During testing, the pre-trained model is frozen to preserve pose priors, and the uncertainty of each joint is utilized to alleviate the overfitting problem.
  • For test time optimization, a projection constraint is used to ensure that the generated results are aligned in the 2D space. Additionally, we design the uncertainty constraint to determine the degree to which each joint can be optimized, which alleviates the overfitting problem.
  • Extensive experiments demonstrate the effectiveness of the UAO framework and its potential applicability for real-world scenarios.
II Related Work
@@ II-A Data-driven Methods
Data-driven approaches involve training a deep neural network to directly regress 3D human poses from provided 2D poses. Early attempts simply use fully-connected networks (FCNs) to lift the 2D keypoint to 3D space [4]. Since the human skeleton can be naturally represented as a graph, various graph convolutional networks (GCNs) based methods [31, 8, 32, 7] have been proposed.
Recently, transformer-based methods [33, 34, 6, 5, 35, 36, 37] have been widely applied to 3D human pose estimation, achieving new breakthroughs. Although these methods can effectively learn pose patterns from datasets, they often exhibit poor generalization ability when faced with unseen data.
II-B Optimization-based Methods
@@
During the early stages, researchers employed pure optimization-based methods [11, 12, 13] for 3D human pose estimation. In the era of deep learning, there are also some methods that integrate optimization strategies. We categorize the existing optimization-based methods into two classes. 1) The first class incorporates a regression-based network with optimization strategies, training the entire structure end-to-end, such as SPIN [18] and PyMAF [38]. 2) The second class follows a post-optimization manner, which is more similar to us, using a pre-trained network to obtain a good initialization, including ISO [20], BOA [17], DynaBOA [16], EFT [19], ESCAPE [21] and CycleAdapt [22]. However, all these previous post-optimization methods typically optimize the network parameters based on specific constraints, which may disrupt the deep prior knowledge obtained from datasets.
Some approaches also directly optimize the pose and shape parameters of parametric human models (e.g., SMPL and SMPL-X), such as SMPLify [39], SMPLify-X [40], KBody [41], and HuMoR [42]. But they are constrained by inaccurate initial values or limited representational space. As a comparison, we set the optimization variable as the predicted 2D pose in latent state, and keep the model parameters frozen during the iterative optimization process. Besides, pose-related constraints are imperative to ensure the appropriate optimization direction. Previous optimization-based methods simply rely on the projection constraint [18, 20, 19]. But due to the depth ambiguity, this imperfect constraint may lead to the wrong optimization direction when over-optimized. To alleviate this problem, we further utilize the uncertainty constraint to ensure that the generated pose lies within the manifold of 3D poses.
II-C Uncertainty-based Methods
Uncertainty learning has attracted considerable attention and has been applied to numerous tasks, such as object detection [43], person re-identification [44], and face recognition [45, 46]. The key idea is that by knowing the reliability of its outputs, a model can optimize its predictions for improved downstream application performance. In the domain of 2D human pose estimation, confidence scores derived from 2D heatmaps are widely used to assess the reliability of each keypoint [47, 39]. For 3D human pose estimation, several studies have explored the use of uncertainty to improve model performance. Han et al. [48] introduced an uncertainty-based framework for 3D human pose estimation to mitigate the negative impacts of joints with high ambiguous. Zhang et al. [49] modeled the depth and the 2D distribution uncertainties separately, optimizing these parameters to achieve accurate 3D poses. In our work, we define keypoint uncertainty as the degree of a keypoint’s deviation from its mean position in 3D space, with a higher uncertainty value indicating a larger allowable divergence radius. Moreover, our approach goes beyond simple uncertainty estimation by incorporating uncertainty constraints into the optimization phase, thereby overcoming the limitations of relying solely on the projection constraint.
III Preliminaries
III-A Theory of Uncertainty Estimation
Assuming the 3D pose follows Gaussian distributions, we reformulate the 2D-to-3D network as a Bayesian Neural Network (BNN) [50], which estimates a Gaussian distribution of the target rather than only predicting the absolute coordinates of points. Specifically, for the 3D joints $\boldsymbol{J}_{i}$ of person $i$ in the dataset, our goal is to estimate the distribution of 3D joints for that individual. During training, the objective is to maximize the likelihood of all observation pairs. Let $\theta$ denote the network‘s learnable parameters. The likelihood function $L(\theta)$ can be formulated as follows:
$L(\theta)=\prod_{i,k}\mathcal{N}(\boldsymbol{J}_{i,k}|\boldsymbol{\mu}_{i,k},\Sigma_{i,k}),$ (1)
where $k$ is the index for the $k$-th joint of a person, $\boldsymbol{\mu}_{i,k}=f_{\mu}(\theta)$ and $\Sigma_{i,k}=f_{\Sigma}(\theta)$ are the mean and covariance matrix of the Gaussian for the $k$-th joint of person $x_{i}$. $f_{\mu}$ and $f_{\Sigma}$ are two network heads predicting joint’s Gaussian mean and covariance. Each joint is considered as a random variable following a normal Gaussian distribution. To further elaborate, the mean $\boldsymbol{\mu}_{i,k}$ captures the most probable location of the $k$-th joint, while the covariance matrix $\Sigma_{i,k}$ characterizes the uncertainty of the distribution around this mean. The likelihood term models the probability of observing the actual 3D joint positions given the estimated means and covariances. In pursuit of the most accurate estimation of the 3D joint positions, we seek the maximum likelihood estimate $\theta^{*}$, which aligns the estimated Gaussian distribution with the observed 3D joint positions, capturing both the most probable locations $\boldsymbol{\mu}_{i,k}$ and the uncertainty $\Sigma_{i,k}$ associated with the $k$-th joint. The formulation of the maximum likelihood function can be achieved by:
$\left.+\ln|\Sigma_{i,k}|\right),$ (2)
where $K$ is the total number of joints per pose. Here we simplify the covariance matrix $\Sigma_{i,k}$ as a diagonal matrix with diagonal elements $\boldsymbol{\sigma}_{i,k}^{2}$, then we get:
III-B Graph Convolutional Network
Graph Convolutional Network (GCN) [51] is capable to capture intricate relationships and structures within graph-structured data. Consider an undirected graph as $G{=}\{V,E\}$, where $V$ is the set of nodes, and $E$ is the set of edges. The edges can be encoded in an adjacency matrix $A{\in}\{0,1\}^{N{\times}N}$. For the input $X_{l}$ of the $l^{th}$ layer, the vanilla graph convolution aggregates the features of the neighboring nodes. The output $X_{l+1}$ of the $l^{th}$ GCN layer can be formulated as:
IV Uncertainty Estimation
@@
For 2D-to-3D lifting, data-driven methods regress the 3D human pose $\hat{\mathbf{J}}^{3D}\in\mathbb{R}^{K\times 3}$ from given 2D pose $\mathbf{J}^{2D}\in\mathbb{R}^{K\times 2}$, where $K$ is the number of skeleton joints. They typically train the model to learn the transformational relationship between 2D pose $\mathbf{J}^{2D}$ and ground truth 3D pose $\mathbf{J}^{3D}$ by minimizing:
$L_{2D}=||\mathbf{J}^{3D}-f(\mathbf{J}^{2D})||_{2},$ (6)
@@
To achieve this, we propose GUMLP, primarily consisting of two parts: a GCN layer and a parallel U-shaped multi-layer perception (UMLP) module, as shown in Figure 2. In contrast to the conventional data-driven model [4, 7], which typically employs only one decoder to make direct predictions, we utilize two distinct decoders to predict $\boldsymbol{\mu}$ and $\boldsymbol{s}$, respectively. The GCN layer focuses on local joint relations and introduces the topological priors of the human skeleton. For the second component, the UMLP module employs a bottleneck structure [53, 8] instead of the original MLP in the transformer [54]. The UMLP module operates in two parallel pathways: Channel UMLP and Spatial UMLP, to capture multi-scale and multi-level features effectively. In the Channel UMLP pathway, the input $\mathbf{X}$ is first processed by a down-projection layer ($\mathbf{X}_{down}$), followed by a middle layer ($\mathbf{X}_{mid}$) that maintains the same dimension, and ultimately, is fed into an up-projection layer ($\mathbf{X}_{up}$). The simplified formulation can be defined as:
@@
[…]
In the Spatial UMLP pathway, given that the number of joints ($K=17$) is already sufficiently sparse, we implement the operations in reverse order along the spatial dimension, starting with $\mathbf{X}_{up}$, followed by $\mathbf{X}_{mid}$, and finally $\mathbf{X}_{up}$. This parallel design along both channel and spatial dimensions allows the UMLP module to extract rich feature representations across different scales and levels.
@@
V Testing-Time Optimization
V-A Optimization Strategy
@@
V-B Projection Constraint
Projection constraint is commonly used in optimization-based methods, such as SPIN [18] and EFT [19]. For the $j$-th 3D pose $\mathbf{J}^{3D}_{j}$, its 2D projection $\mathbf{J}^{2D}_{j}$ should conform to the principles of perspective projection [13], which can be expressed as follow:
$\mathbf{J}^{3D}_{j}\cdot\mathbf{P}=\mathbf{J}^{2D}_{j},$ (9)
where the projection matrix $\mathbf{P}\in\mathbb{R}^{3\times 2}$ contains the intrinsic parameters of the camera. In real scenarios, the intrinsic parameters of the camera can be obtained from camera specifications or can be inferred solely from the input images [57]. During the testing-time optimization, the 2D projection $\hat{\mathbf{J}}^{2D}_{j}=\hat{\mathbf{J}}^{3D}_{j}\cdot\mathbf{P}$ of the predicted 3D pose $\hat{\mathbf{J}}^{3D}_{j}$ is expected to closely align with the given 2D pose $\mathbf{J}^{2D}_{j}$. A large error between them indicates that the estimated 3D pose $\hat{\mathbf{J}}^{3D}_{j}$ may be problematic and needs to be corrected. The formulation of projection constraint $L^{opt}_{P,j}$ can be expressed as:
$L^{opt}_{P,j}=||\hat{\mathbf{J}}^{2D}_{j}-\mathbf{J}^{2D}_{j}||_{2}.$ (10)
The purpose of the projection constraint is to enforce the relationship between the estimated 3D pose $\hat{\mathbf{J}}^{3D}_{j}$ and its corresponding 2D projection $\hat{\mathbf{J}}^{2D}_{j}$ to follow the geometric principles of perspective projection, thereby improving the accuracy of the pose estimation.
Input: 2D pose $\mathbf{J}_{j}^{2D}$, pre-trained model $f^{*}$ with fixed parameters, projection matrix $\mathbf{P}$, optimization steps $T$
Output: 3D pose $\hat{\mathbf{J}}_{j}^{3D}$
1 Initialization: $\mathbf{z}_{j}=\mathbf{J}_{j}^{2D}$, $iter=0$
2 $\hat{\boldsymbol{\mu}}_{j}=f_{\boldsymbol{\mu}}^{*}(\mathbf{J}_{j}^{2D}),\hat{\boldsymbol{s}}_{j}=f_{\boldsymbol{s}}^{*}(\mathbf{J}_{j}^{2D})$
3 while  $iter<T$ do
4 $\hat{\mathbf{J}}_{j}^{3D}=f_{\boldsymbol{\mu}}^{*}(\mathbf{z}_{j})$
5 $L_{P,j}^{opt}=\|\hat{\mathbf{J}}_{j}^{3D}\cdot\mathbf{P}-\mathbf{J}_{j}^{2D}\|_{2}$
6 $L_{U,j}^{opt}=\left\|\hat{\boldsymbol{\mu}}_{j}-\hat{\mathbf{J}}_{j}^{3D}\right\|^{2}/2\exp(\hat{\boldsymbol{s}}_{j})$
7 $L_{total,j}^{opt}=\lambda_{P}\cdot L_{P,j}^{opt}+\lambda_{U}\cdot L_{U,j}^{opt}$
8 $\mathbf{z}_{j}\leftarrow Adam(\mathbf{z}_{j},\nabla L_{total,j}^{opt})$;
9 $iter=iter+1$
10
[…]
return $\hat{\mathbf{J}}_{j}^{3D}$